Find the value of sin (A+B) when tan(A+B)=√3 now find solution

Grade:10

1 Answers

Arun
6 years ago
Dear student
 
tan (A+B) = sqrt (3)
tan (A + B) = tan 60°
 
Hence
 
A+B = 60 °
 
Now
sin (A+ B) = sin 60° = (1/2) * sqrt(3)
 
 
Regards
Arun (askIITians forum expert)